OpenStudy (anonymous):

In a certain candy store, 3 pounds of candy and 2 pounds of mints cost $10.80, and 1 pound of candy and 3 pounds of mints cost $5.35. What is the cost per pound of the mints? mints are 65¢ per lb mints are 75¢ per lb mints are 85¢ per lb

OpenStudy (anonymous):

it has to be a large number so im thinking C

OpenStudy (anonymous):

well, first you have to figure out how much the candy itself is per pound. then you pick the 1st example you have, and subtract the amount you have by 3. then you divide what you have by 2(cuz thats how many pounds u have in that example) and you have ur answer

OpenStudy (anonymous):

how do i gigure out the first part you jsut said

OpenStudy (anonymous):

figure just***

OpenStudy (anonymous):

you take the # they give you all together(10.80) then uou divide that by 3

OpenStudy (anonymous):

you***

OpenStudy (anonymous):

kk....one sec

OpenStudy (anonymous):

i got 3.6 is that correct

OpenStudy (anonymous):

i think I did that wrong...hold on, im srry

OpenStudy (anonymous):

npnp

OpenStudy (anonymous):

ok... Let c and m be the cost, $ per lb, of candy and mints respectively. Solve the following equations for c and m. {3 c + 2 m = 10.80, 1 c + 3 m = 5.35} {c = $3 .10/lb, m = $0 .75/lb} mints are 75¢ per lb

OpenStudy (anonymous):

do u get it now?

OpenStudy (anonymous):

ur awesome turtleluver

OpenStudy (anonymous):

thx?